Required Information (The following information applies to the questions displayed below] Delph Company uses a job-order costing system and has two manufacturing departments-Molding and Fabrication. The company provided the following estimates at the beginning of the year Fabrication Molding 31,00e $ 700,00e 5.30 Total Machine-hours Fixed manufacturing overhead costs Variable manufacturing overhead cost per machine-hour 41,800 72,eee $ 920, eee $ 220,0e0 5.30 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs- Job D-70 and Job C-200. It provided the following information related to those two jobs: Job D-70: Direct materials cost Direct labor cost Machine-hours Molding Fabrication S 375,eee $ 322,000 $697,000 $ 250,eee $ 170,eee $ 428,eee 25,eee Total 6,eee 31,eee Job C-200: Direct materials cost Direct labor cost Machine-hours Molding Fabrication Total $ 230,eee $ 270,eee$ 5ee,eee $ 170,eee $ 220,eee $ 390,eee 6,eee 35, eee 41,e00 Delph had no underapplied or overapplied manufacturing overhead during the year 2 Assume Delph uses departmental predetermined overhead rates based on machine-hours. a. Compute the departmental predetermined overhead rates. b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200 c. If Delph establishes bid prices that are 140% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200? d. What is Delph's cost of goods sold for the year? Complete the question by entering your answers in the tabs given below. Required 2A Required 2B Required 2C Required 2D Compute the departmental predetermined overhead rates. (Round the final answers to 2 decimal places.) Pre determined overhead rates Molding Department per MH Fabrication Department per MH
Required 2B Required 2C Required 2D Required 2A Compute the total manufacturing cost assigned to Job D-70 and Job C-200. (Round your intermediate decimal places. Round your final answers to nearest whole dollar amount.) Total manufacturing cost Job D-70 Job C-200
Required Required 2A Required 2B Required 2D If Delph establishes bid prices that are 140% of total manufacturing costs, what bid prices would it D-70 and Job C-200? (Round your intermediate calculations to 2 decimal places. Round your final a dollar amount.) Bid price Job D-70 Job C-200 Required 2B Required 2D >
Required 2A Required 2B Required 2C Required 2D What is Delph's cost of goods sold for the year? (Round your in answer to nearest whole dollar amount.) Cost of goods sold
Answer #1

1) Departmental overhead rate :

Molding = (700000/31000)+5.30 = 27.88 per MH

painting = (220000/41000)+5.30 = 10.67 per MH

2) calculate total manufacturing cost

Job D-70 Job C-200
Direct material 697000 500000
Direct labor 420000 390000
Overhead 31000*27.88 = 864280 41000*10.67 = 437470
Total manufacturing cost 1981280 1327470

3) Bid price :

Job D-70 = 1981280*1.4 = 2773792

Job C-200 = 1327470*1.4 = 1858458

4) Cost of goods sold = 1981280+1327470 = 3308750